OpenStudy (anonymous):

∫[e^(3t)sin(3t),dt]= ;limit(0,π)

OpenStudy (anonymous):

Alma, you have to apply integration by parts twice to this integral. Set u=e^(3t) and dv=sin(3t)dt to start.